OpenStudy (surryyy):

Given f(x) = x + 11 and g(x) = x - 7, what is f(x) + g(x)? -77x 18x 2x + 4 -2 - 77x

OpenStudy (camerondoherty):

Just add c: (x + 11) + (x - 7) First we add the common terms. x + x = ?

OpenStudy (surryyy):

would it be 2x+4?

OpenStudy (camerondoherty):

Yep! because x + x = 2x and 11 - 7 = 4 And the final result would be 2x + 4 Great Job! c:

OpenStudy (surryyy):

thnx.

OpenStudy (camerondoherty):

Anytime c: